Problem 6-49 Weighted-Average Method; Transferred-in Costs; Spoilage (Appendix) (LO 6-2, 6-4, 6-6, 6- 9) APCO Company manufac3. Compute the cost per equivalent unit for the fully assembled bike. 4. Compute the amount of the total production cost of $Required 1 Required 2 Required 3 Required 4 Compute the equivalent units of production for the year for a. Bicycles transferrRequired 1 Required 2 Required 3 Required 4 Compute the cost per equivalent unit for the fully assembled bike. decimal placesRequired 1 Required 2 Required 3 Required 4 Compute the amount of the total production cost of $1,708,650 that will be associ
Answer #1

Part 1

Normal spoilage = 2689 units

Abnormal spoilage = 6701 units

Part 2

Transferred-in

Materials

Conversion

Equivalent units of production

59110

57000

52917

Part 3

Cost per equivalent unit

$29.49

Part 4

Normal spoilage

164171

Goods units completed in the assembly department

1031485

Abnormal spoilage

409115

Ending work in process inventory

103879

Explanation

Process Cost Report

Inspection Point

70%

Percent complete

Whole units

Materials

Conversion costs

Transferred-in

Beginning WIP

3110

Units started or Trans-in

56000

Total to account for

59110

Units Finshed or Trans-out

45500

100%

100%

100%

Normal spoilage

2689

100%

70%

100%

Abnormal spoilage

6701

100%

70%

100%

Ending WIP

4220

50%

20%

100%

Total accounted for

59110

Supporting Calculation: Determining Number of Spoiled Units (bikes)

Normal spoilage (in units):

Bikes passing through assembly dept during May

59110

Less: bikes not inspected during current year:

Beginning WIP inventory

3110

End WIP inventory

4220

5330

Bikes that reached the insp point this year

53780

Normal defect (spoilage) rate

5%

Normal spoilage (units)

2689

Abnormal spoilage (defective units):

Total bikes lost

9390

(59110-45500-4220)

Normal spoilage, in units

2689

Amount of abnormal spoilage, in units

6701

Transferred In

59110

57000

52917

Materials

Conversion

Calculating Equivalent Unit Costs:

Added by Assembly Department

Trans-in

Materials

Conversion

Total

Beginning WIP

93200

8310

13800

115310

Current Costs

1246600

102120

244620

1593340

Total

1339800

110430

258420

1708650

Cost per EU

22.67

1.94

4.88

29.49

COST SUMMARY - WEIGHTED AVERAGE

Finished Goods

45500

Good production

1031485

Normal spoilage

164171

TOTAL

1195656

Abnormal Spoilage

409115

Ending WIP

4220

Trans-in

95667

Materials

4093

Conversion

4119

103879

Total Costs Accounted For

1708650

Total spoilage = 1708650-(1031485+103879) = 573286

Normal spoilage = 573286*2689/9390 = 164171

Abnormal spoilage = 573286*6701/9390 = 409115